A speeder tries to explain to the police that the yellow warning lights she was approaching on the side of the road looked green to her because of the Doppler shift. How fast would she have been traveling if yellow light of wavelength 577.1 nm had been shifted to green with a wavelength of 563.4 nm? Note: For speeds less than 0.03c, the equation below will lead to a value for the observed frequency accurate to approximately two significant digits.
Answer #1


yellow light wavelength Ly=577.1 nm

green light wavelength Lg=563.4 nm

let speed is u

use,

frequency observed is f o=fs*(1+u/c)

c/Lg=C/Ly*(1+u/c)

Ly/Lg=(1+u/c)

u=c*(Ly/Lg-1)

u=3*10^8*((577.1/563.4 -1)

u=7.29*10^6 m/sec